# Sandwiched singularities and nearly Lefschetz fibrations

### v2

@@ Abstract.
We study Milnor fibers and symplectic fillings of links of sandwiched singularities, with the goal of contrasting their algebro-geometric deformation theory and symplectic topology. In the algebro-geometric setting, smoothings of sandwiched singularities are described by de Jong–van Straten’s theory: all Milnor fibers are generated from deformations of a singular plane curve germ associated to the surface singularity. We develop an analog of this theory in the symplectic setting, showing that all minimal symplectic fillings of the links are generated by certain immersed disk arrangements resembling de Jong–van Straten’s picture deformations. This paper continues our previous work for a special subclass of singularities; the general case has additional difficulties and new features. The key new ingredient in the present paper is given by spinal open books and nearly Lefschetz fibrations: we use recent work of Roy–Min–Wang to understand symplectic fillings and encode them via multisections of certain Lefschetz fibrations. As an application, we discuss arrangements that generate unexpected Stein fillings that are different from all Milnor fibers, showing that the links of a large class of sandwiched singularities admit unexpected fillings.
@@ 1. Introduction
In this paper, we compare and contrast deformation theory and symplectic topology of certain rational surface singularities. Extending our work in [PS23], we study Stein fillings of the link of a singularity and compare them to Milnor fibers of possible smoothings. Let $X\subset\mathbb{C}^{N}$ be a singular complex surface with an isolated singularity at the origin. The intersection $Y=X\cap S^{2N-1}_{r}$ with the sphere of a small radius $r>0$ is a smooth 3-manifold called the link of the singularity $(X,0)$. The canonical contact structure $\xi$ on $Y$ is the distribution of complex tangencies to $Y$; the contact link $(Y,\xi)$ is independent of choices, up to contactomorphism. If $(X,0)$ admits a smoothing (that is, a deformation to a smooth surface, called a Milnor fiber), then the Milnor fiber carries a Stein structure and gives a Stein filling of $(Y,\xi)$. For a general surface singularity, the collection of Stein fillings is typically much larger than that of Milnor fibers, [AO14, AO18, BNP25]. However, restrictions on the topology of the link often leads to rigidity phenomena, where it turns out that all possible Stein fillings are generated by Milnor fibers. For example, when the link is a lens space, all minimal symplectic fillings are given by the Milnor fibers of the corresponding cyclic quotient singularity, [Lis08, NPP10b]; analogous statements hold for certain other classes of singularities, [OO03, BO12, PPSU18]. We will focus on sandwiched singularities, a subclass of rational singularities that can be defined by a combinatorial hypothesis on the dual resolution graph. The links of rational singularities are L-spaces in the Heegaard Floer sense [Ném08]; this, together with other properties, significantly obstructs symplectic fillings. We will see that for sandwiched singularities, the symplectic topology of the fillings is still largely informed by the algebro-geometric constructions but rigidity begins to break down. This is the most interesting borderline case.
@@
Unexpected fillings uncover the difference between the algebraic and symplectic settings. These are Stein domains that look quite similar to but still different from the “standard” fillings provided by Milnor fibers. As such, these constructions may have potential applications in 4-manifold topology; one could try to use unexpected fillings to produce interesting closed 4-manifolds. It would also be interesting to find other ways to construct further unexpected fillings.
The paper is organized as follows. Section 2 contains the main definitions and statements related to sandwiched singularities and the De Jong–Van Straten theory. We define DJVS immersed disk arrangements and state a precise version of Theorem 1.1. Section 3 discusses spinal open books and nearly Lefschetz fibrations and explains how they arise in our setting. We show that the contact link of a sandwiched singularity is supported by a planar spinal open book, and therefore each Stein filling has the structure of a nearly Lefschetz fibration. In Section 4, we develop a dictionary for translating between DJVS disk arrangements and nearly Lefschetz fibrations, and finish the proof of Theorem 1.1. Section 5 explains how to distinguish the fillings using the incidence data; we summarize our previous results on unexpected fillings and prove Theorem 1.2. We have made an effort to include sufficient background and details to make the paper reasonably self-contained.
Acknowledgements: We are grateful to Márton Beke, Andras Némethi, and Mark Powell for helpful conversations, and especially to Luya Wang for explaining spinal open books to us. Some of the initial discussions motivating this work took place at the conference “New structures in low-dimensional topology” in Budapest, July 2024 and at the workshop “What’s your trick?” in Banff, August 2024. The authors discussed the final stages of this project during the Georgia Topology Conference in May 2025. OP has been partially supported by the NSF grant DMS 2304080. LS has been partially supported by NSF CAREER grant DMS 2042345 and Sloan grant FG 2021-16254.
@@ 2.1. Decorated germs and picture deformations
Figure 1. A resolution graph of a sandwiched singularity together with an augmentation with 1-1 curves, decorated by transverse disks. Below is the sequence of blow-downs and the image of the disks under these blow-downs. In this case, the resulting curve configuration consists of two simple cusps intersecting each other with multiplicity $7$. By tracking the multiplicities of the exceptional spheres at each stage, one can see that the weight $w_{i}$ of each component is $8$.
@@ 3.2. Nearly Lefschetz fibrations
A nearly Lefschetz fibration induces a spinal open book on its boundary. The paper part of the open book is given by the fiber bundle $d_{vert}W\to\partial D^{2}$, so the page of the open book is the general fiber of the nearly Lefschetz fibration. The spine is given by $\partial_{hor}W$, which is the union of solid tori. The meridional disk of each solid torus component (“vertebra” in the terminology of [MRW24, LHMW18]) is a branched covering of the base $D^{2}$ of the fibration, with simple branch points that correspond to exotic fibers. Since the boundary of this meridional disk is a $d_{i}$-fold cover of $\partial D^{2}$, where $d_{i}$ is the multiplicity of the corresponding binding component in the spinal open book, the number of branch points is determined by the Riemann–Hurwitz formula. It follows that the number of exotic fibers in a nearly Lefschetz fibration with a given spinal open book boundary is determined by the open book and will be the same for all such fibrations.
Nearly Lefschetz fibrations naturally arise on the complements of positive multisections of Lefschetz fibrations, [BH16]. A positive multisection is a simple branched cover of the base, with the branch points modeled on the projection to $\mathbb{C}_{x}$ of the curve $y^{2}=x$ in $\mathbb{C}^{2}=\mathbb{C}_{x}\times\mathbb{C}_{y}$. A generic projection of a smooth complex curve in $\mathbb{C}^{2}$ gives a multisection of the product fibration $\mathbb{C}^{2}\to\mathbb{C}$. Baykur and Hayano [BH16] show that given a positive multisection in a positive allowable Lefschetz fibration, the complement of the multisection has the structure of positive allowable nearly Lefschetz fibration. The fiber of this new fibration is the fiber of the original fibration with holes, and the new monodromy is obtained as the composition of the original monodromy with the boundary interchanges that correspond to branch points. See [BH16, Theorem 1.1] and also [MRW24, Theorem 2.13] for a statement that is geared more closely to our case.
@@
Using the local model calculation of [PS23, Lemma 3.2] in a neighborhood of an intersection point of $\Gamma_{i}$’s, we see that the blow-up at the marked point $p_{j}$ produces a Lefschetz fibration whose singular fiber contains the exceptional sphere. The map is the composition of the blow-down with the projection to $D_{x}$. The strict transforms of the disks $\Gamma_{i}$ meeting at the point $p_{j}$ pass through distinct points of that sphere since the interesection is transverse. The corresponding vanishing cycle in the nearby fiber of this Lefschetz fibration encloses the intersections of these disks $\Gamma_{i}$ with the fiber. Thus, after deleting the neighborhood of the strict transforms of the $\Gamma_{i}$, these vanishing cycles become essential curves in the fiber, so the Lefschetz singularities will be allowable. The strict transforms $\tilde{\Gamma}_{i}$ form multisections of the Lefschetz fibration on $(D_{x}\times D_{y})\#_{n}\overline{\mathbb{C}\textup{P}}\,\!^{2}$, so by the arguments of [MRW24, BH16] the local model for the complement of (the tubular neighborhoods of) $\tilde{\Gamma}_{i}$ near a branch point agrees with the local model of an exotic fiber. (Note that Baykur–Hayano’s theorem [BH16] does not apply as stated because the original fibration on the blow-up of the disk is not allowable, but in our setting the argument goes through.) Thus, $W_{(\Gamma,p)}$ carries a positive allowable nearly Lefschetz fibration with the exotic fibers corresponding to the branch points. ∎
@@ 4.2. Braided wiring diagrams:
We would like to encode a DJVS disk arrangement $\Gamma=\cup\Gamma_{i}$ in $D_{x}\times D_{y}$. The branch points of the projection arise when $\Gamma_{i}$ is tangent to a fiber of $\pi_{x}$, so we refer to these critical points of the projection as tangencies. The key information is the branch points, the intersections, and how these connect to each other. We will show that $\Gamma$ can be determined up to isotopy by a 1-dimensional singular braid with the tangencies and intersection points specified, called a braided wiring diagram. We generalize the notion of braided wiring diagram (first used by Arvola [Arv92] to study complex line arrangements) to encode tangencies. Combinatorially, a braided wiring diagram can be encoded by a sequence $(b_{0},S_{1},b_{1},S_{2},\dots,b_{N-1},S_{N},b_{N})$ where $b_{i}$ is a braid word on $n$ strands, and $S_{i}$ is a singularity–either a transverse intersection between a specified collection of adjacent strands or a tangency between two specified adjacent strands. (One could easily generalize this to allow more singularities, as long as a model is specified for each singularity type.) See Figure 6 for an example.
Figure 6. Braided wiring diagram $(1,T_{2,3},\sigma_{1}^{-1}\sigma_{3}^{-1},T_{2,3},1,I_{1,2},1,I_{1,2},1,I_{1,2},\sigma_{3}^{-1},I_{2,3},$ $\sigma_{2}^{-1},I_{1,2},\sigma_{3}\sigma_{2},I_{3,4},1,I_{1,2,3})$. Strands are enumerated from bottom to top with local indices (i.e. height at that position in the diagram). The singularities are labeled $T_{i,i+1}$ for a tangency between the $i^{th}$ and $i+1^{st}$ strands, and $I_{i,...,i+k}$ for a transverse intersection between the $i^{th}$ through $(i+k)^{th}$ strands. Braid words are read right to left with multiplicative notation, $\sigma_{i}$ indicating a crossing between the $i^{th}$ and $(i+1)^{st}$ strands where the $i^{th}$ strand crosses under the $(i+1)^{st}$ strand from right to left.
@@
Fix $0<\mu\ll\pi/2$. Our model for a tangency will be the complex curve $\{(x,y)\in\mathbb{C}^{2}\mid e^{i\mu}x=y^{2}\}$ with the tangency of $\pi_{x}$ at the origin. The braided wiring diagram and braid monodromy near this model are shown in Figure 7. The preimage of the model over the interval in the $Re(x)$ axis includes the singularity, and gives the picture for the braided wiring diagram near a tangency. The braid monodromy about a small circle oriented counterclockwise around the singularity is a single positive half twist on two strands. We break up this circle into two half-circles showing the push-off of the singular braid in the positive and negative imaginary directions in $\mathbb{C}_{x}$. Note that the positive push-off given by the top green braid is oriented right to left under the counterclockwise orientation on the circle, which makes the projection of positive crossing in Figure 7 (rotating counterclockwise in the oriented $\mathbb{C}_{y}$ while moving from right to left) look different from some other knot theory conventions. By taking the inverse of the braid corresponding to the positive push-off and composing with the braid corresponding to the negative push-off, we get the braid monodromy around the circle as a standard positive half twist on two strands.
@@
The model for a transverse intersection is given by a pencil of complex lines which are the complexification of real transverse lines. See [PS23, CS97] for more details on braided wiring diagrams with only transverse intersection singularities and models. Figure 8 shows a braided wiring diagram including both tangency and transverse intersection models, along with the pushoffs in the positive and negative imaginary directions.
@@
Given a DJVS arrangement $(\Gamma,\{p_{j}\})$, Theorem 3.8 gives a construction of an associated manifold $W$ with a nearly Lefschetz fibration. We have seen that the arrangement can be encoded in a braided wiring diagram (with marked points). A nearly Lefschetz fibration can be encoded by an ordered list of vanishing cycles (specifying Lefschetz singularities) and vanishing arcs (specifying exotic fibers). Our next goal is to explain how to translate back and forth between the braided wiring diagram and the vanishing cycles and vanishing arcs resulting from Theorem 3.8.
@@
Let $(\Gamma,\{p_{j}\})$ be a DJVS arrangement with marked points and the corresponding braided wiring diagram $(b_{0},S_{1},b_{1},S_{2},\dots,b_{N-1},S_{N},b_{N})$, where a free marked count will also be considered a singular point $S_{i}$, treated as an intersection point with only one strand. Fix an identification of the planar fiber $P$, a disk with disjoint holes of equal radius centered along a straight line. If the singularity $S_{i}$ is an intersection of strands $\{j,j+1,\dots,j+k\}$, let $A_{i}$ denote a convex curve containing holes $j,j+1,\dots,j+k$, and let $\phi_{i}$ denote the positive half twist $\Delta_{j,j+k}$ on strands $j,j+1,\dots,j+k$ (whose square is the positive Dehn twist on $A_{i}$). In the case of a free marked point where $k=0$, $A_{i}$ is the boundary parallel curve around the $j^{th}$ hole, $\Delta_{j,j}$ is the positive half twist on $A_{i}$ (whose square is a positive Dehn twist on $A_{i}$), and $\phi_{i}=\Delta_{j,j}$. If the singularity $S_{i}$ is a tangency between strands $j,j+1$, let $A_{i}$ denote a straight line arc connecting hole $j$ and $j+1$, and let $\phi_{i}=id$. Let $(V_{1},V_{2},\dots,V_{N})$ be an ordered collection of vanishing cycles and arcs corresponding to the nearly Lefschetz fibration on the manifold $W_{(\Gamma,p)}$ produced by Theorem 3.8. Then up to Hurwitz equivalence,
@@
Recall that $W_{(\Gamma,p)}=[(D_{x}\times D_{y})\#_{n}\overline{\mathbb{C}\textup{P}}\,\!^{2}]\setminus\cup\nu(\widetilde{\Gamma}_{i})$, and the Lefschetz fibration is the composition of the blow-down with the projection to $D_{x}$. Let $\alpha:I\to D_{x}$ be the arc over which lies the braided wiring diagram with data $(b_{0},S_{1},b_{1},S_{2},\dots,b_{N-1},S_{N},b_{N})$. We will define vanishing cycles and arcs in the fiber over $\alpha(0)$. We can identify the vanishing cycles by choosing an ordered collection of loops, one about each singular value, such that the composition of the loops is homotopic to an embedded loop convexly surrounding all the singular values. The monodromy about each loop will be a positive (right handed) Dehn twist about the vanishing cycle or a positive boundary interchange about the vanishing arc, depending on whether the singularity is Lefschetz or exotic respectively. Thus if we can choose such a collection of loops in $D_{x}$, determine the monodromy above each loop, and realize this monodromy as a positive Dehn twist about a curve, then we will have identified a set of vanishing cycles for the Lefschetz fibration. Other collections of vanishing cycles (corresponding to different choices of loops or base fiber) will be related to this set via Hurwitz equivalence (which includes both Hurwitz moves and global conjugation).
Let $t_{1},\dots,t_{N}\in(0,1)$ be such that $\alpha(t_{i})$ is the image of the singularity $S_{i}$. Let $t_{0}=0$, $t_{N+1}=1$. Choose a sufficiently small $\delta>0$ such that $\alpha|{[t_{i}-\delta,t_{i}+\delta]}(t)=(t_{i}-t)+q_{i}$ is parallel to the real axis in $D_{x}$ and the braids $b_{i}$ lie over the intervals $\alpha([t_{i}+\delta,t_{i+1}-\delta])$ for $i=0,\dots,N$. Let $h_{i}(t)=\delta e^{-it/\pi}+q_{i}$ be the path along the lower semi-circle of radius $\delta$ around $q_{i}$. Let $g_{i}(t)=\delta e^{it/(2\pi)}+q_{i}$ be the counterclockwise circle of radius $\delta$ around $q_{i}$. Then the loop $l_{i}$ is the concatenation of the following paths:

### v3

@@ Abstract.
We study Milnor fibers and symplectic fillings of links of sandwiched singularities, with the goal of contrasting their algebro-geometric deformation theory and symplectic topology. In the algebro-geometric setting, smoothings of sandwiched singularities are described by de Jong–van Straten’s theory: all Milnor fibers are generated from deformations of a singular plane curve germ associated to the surface singularity. We develop an analog of this theory in the symplectic setting, showing that all minimal symplectic fillings of the links are generated by certain immersed disk arrangements resembling de Jong–van Straten’s picture deformations. This paper continues our previous work for a special subclass of singularities; the general case has additional difficulties and new features. The key new ingredient in the present paper is given by spinal open books and nearly Lefschetz fibrations: we use recent work of Min–Roy–Wang to understand symplectic fillings and encode them via multisections of certain Lefschetz fibrations. As an application, we discuss arrangements that generate unexpected Stein fillings that are different from all Milnor fibers, showing that the links of a large class of sandwiched singularities admit unexpected fillings.
@@ 1. Introduction
In this paper, we compare and contrast deformation theory and symplectic topology of certain rational surface singularities. Extending our work in [PS23], we study Stein fillings of the link of a singularity and compare them to Milnor fibers of possible smoothings. Let $X\subset\mathbb{C}^{N}$ be a singular complex surface with an isolated singularity at the origin. The intersection $Y=X\cap S^{2N-1}_{r}$ with the sphere of a small radius $r>0$ is a smooth 3-manifold called the link of the singularity $(X,0)$. The canonical contact structure $\xi$ on $Y$ is the distribution of complex tangencies to $Y$; the contact link $(Y,\xi)$ is independent of choices, up to contactomorphism. If $(X,0)$ admits a smoothing (that is, a deformation to a smooth surface, called a Milnor fiber), then the Milnor fiber carries a Stein structure and gives a Stein filling of $(Y,\xi)$. For a general surface singularity, the collection of Stein fillings is typically much larger than that of Milnor fibers, [AO14, AO18, BNP24]. However, restrictions on the topology of the link often leads to rigidity phenomena, where it turns out that all possible Stein fillings are generated by Milnor fibers. For example, when the link is a lens space, all minimal symplectic fillings are given by the Milnor fibers of the corresponding cyclic quotient singularity, [Lis08, NPP10b]; analogous statements hold for certain other classes of singularities, [OO03, BO12, PPSU18]. We will focus on sandwiched singularities, a subclass of rational singularities that can be defined by a combinatorial hypothesis on the dual resolution graph. The links of rational singularities are L-spaces in the Heegaard Floer sense [Ném08]; this, together with other properties, significantly obstructs symplectic fillings. We will see that for sandwiched singularities, the symplectic topology of the fillings is still largely informed by the algebro-geometric constructions but rigidity begins to break down. This is the most interesting borderline case.
@@
Unexpected fillings uncover the difference between the algebraic and symplectic settings. These are Stein domains that look quite similar to but still different from the “standard” fillings provided by Milnor fibers. As such, these constructions may have potential applications in 4-manifold topology; one could try to use unexpected fillings to produce interesting closed 4-manifolds. For example, in our next paper [BPS25] joint with Márton Beke, we were able to construct rational homology disk Stein fillings for certain links of singularities where rational homology disk smoothings do not exist. It would be interesting to find other ways to construct further unexpected fillings and their further applications.
The paper is organized as follows. Section 2 contains the main definitions and statements related to sandwiched singularities and the de Jong–van Straten theory. We define DJVS immersed disk arrangements and state a precise version of Theorem 1.1. Section 3 discusses spinal open books and nearly Lefschetz fibrations and explains how they arise in our setting. We show that the contact link of a sandwiched singularity is supported by a planar spinal open book, and therefore each Stein filling has the structure of a nearly Lefschetz fibration. In Section 4, we develop a dictionary for translating between DJVS disk arrangements and nearly Lefschetz fibrations, and finish the proof of Theorem 1.1. Section 5 explains how to distinguish the fillings using the incidence data; we summarize our previous results on unexpected fillings, prove Theorem 1.2, and include a discussion of deformation (non)realizability of arrangements. We have made an effort to include sufficient background and details to make the paper reasonably self-contained.
Acknowledgements: We are grateful to Márton Beke, Andras Némethi, and Mark Powell for helpful conversations, and especially to Luya Wang for explaining spinal open books to us. We thank Stepan Orevkov for very helpful discussions of deformation realizability of curve arrangements. Some of the initial discussions motivating this work took place at the conference “New structures in low-dimensional topology” in Budapest, July 2024 and at the workshop “What’s your trick?” in Banff, August 2024. The authors discussed the final stages of this project during the Georgia Topology Conference in May 2025. OP has been partially supported by the NSF grant DMS 2304080. LS has been partially supported by NSF CAREER grant DMS 2042345 and Sloan grant FG 2021-16254.
@@ 2.1. Decorated germs and picture deformations
Figure 1. A resolution graph of a sandwiched singularity together with an augmentation with 1-1 curves, decorated by transverse disks. Below is the sequence of blow-downs and the image of the disks under these blow-downs. In this case, the resulting curve configuration consists of two simple cusps intersecting each other with multiplicity $7$. By tracking the multiplicities of the exceptional spheres at each stage, one can see that the weight $w_{i}$ of each component is $6$.
@@ 3.2. Nearly Lefschetz fibrations
A nearly Lefschetz fibration induces a spinal open book on its boundary. The paper part of the open book is given by the fiber bundle $\partial_{vert}W\to\partial D^{2}$, so the page of the open book is the general fiber of the nearly Lefschetz fibration. The spine is given by $\partial_{hor}W$, which is the union of solid tori. The meridional disk of each solid torus component (“vertebra” in the terminology of [MRW24, LHMW18]) is a branched covering of the base $D^{2}$ of the fibration, with simple branch points that correspond to exotic fibers. Since the boundary of this meridional disk is a $d_{i}$-fold cover of $\partial D^{2}$, where $d_{i}$ is the multiplicity of the corresponding binding component in the spinal open book, the number of branch points is determined by the Riemann–Hurwitz formula. It follows that the number of exotic fibers in a nearly Lefschetz fibration with a given spinal open book boundary is determined by the open book and will be the same for all such fibrations.
Nearly Lefschetz fibrations naturally arise on the complements of positive multisections of Lefschetz fibrations, [BH16b]. A positive multisection is a simple branched cover of the base, with the branch points modeled on the projection to $\mathbb{C}_{x}$ of the curve $y^{2}=x$ in $\mathbb{C}^{2}=\mathbb{C}_{x}\times\mathbb{C}_{y}$. A generic projection of a smooth complex curve in $\mathbb{C}^{2}$ gives a multisection of the product fibration $\mathbb{C}^{2}\to\mathbb{C}$. Baykur and Hayano [BH16b] show that given a positive multisection in a positive allowable Lefschetz fibration, the complement of the multisection has the structure of positive allowable nearly Lefschetz fibration. The fiber of this new fibration is the fiber of the original fibration with holes, and the new monodromy is obtained as the composition of the original monodromy with the boundary interchanges that correspond to branch points. See [BH16b, Theorem 1.1] and also [MRW24, Theorem 2.13] for a statement that is geared more closely to our case.
@@
Using the local model calculation of [PS23, Lemma 3.2] in a neighborhood of an intersection point of $\Gamma_{i}$’s, we see that the blow-up at the marked point $p_{j}$ produces a Lefschetz fibration whose singular fiber contains the exceptional sphere. The map is the composition of the blow-down with the projection to $D_{x}$. The strict transforms of the disks $\Gamma_{i}$ meeting at the point $p_{j}$ pass through distinct points of that sphere since the interesection is transverse. The corresponding vanishing cycle in the nearby fiber of this Lefschetz fibration encloses the intersections of these disks $\Gamma_{i}$ with the fiber. Thus, after deleting the neighborhood of the strict transforms of the $\Gamma_{i}$, these vanishing cycles become essential curves in the fiber, so the Lefschetz singularities will be allowable. The strict transforms $\tilde{\Gamma}_{i}$ form multisections of the Lefschetz fibration on $(D_{x}\times D_{y})\#_{n}\overline{\mathbb{C}\textup{P}}\,\!^{2}$, so by the arguments of [MRW24, BH16b] the local model for the complement of (the tubular neighborhoods of) $\tilde{\Gamma}_{i}$ near a branch point agrees with the local model of an exotic fiber. (Note that Baykur–Hayano’s theorem [BH16b] does not apply as stated because the original fibration on the blow-up of the disk is not allowable, but in our setting the argument goes through.) Thus, $W_{(\Gamma,p)}$ carries a positive allowable nearly Lefschetz fibration with the exotic fibers corresponding to the branch points. ∎
@@ 4.2. Braided wiring diagrams:
We would like to encode a DJVS disk arrangement $\Gamma=\cup\Gamma_{i}$ in $D_{x}\times D_{y}$. The branch points of the projection arise when $\Gamma_{i}$ is tangent to a fiber of $\pi_{x}$, so we refer to these critical points of the projection as tangencies. The key information is the branch points, the intersections, and how these connect to each other. We will show that $\Gamma$ can be determined up to isotopy by a 1-dimensional singular braid with the tangencies and intersection points specified, called a braided wiring diagram. We generalize the notion of braided wiring diagram (first used by Arvola [Arv92] to study complex line arrangements) to encode tangencies. Combinatorially, a braided wiring diagram can be encoded by a sequence $(b_{0},S_{1},b_{1},S_{2},\dots,b_{N-1},S_{N},b_{N})$ where $b_{i}$ is a braid word on $n$ strands, and $S_{i}$ is a singularity–either a transverse intersection between a specified collection of adjacent strands or a tangency between two specified adjacent strands. (One could easily generalize this to allow more singularities, as long as a model is specified for each singularity type.) For brevity, we will also write a sequence of the form $(b_{0},S_{1},b_{1},S_{2},\dots,b_{N-1},S_{N},b_{N})$ as a word $b_{0}S_{1}b_{1}S_{2}\dots b_{N-1}S_{N}b_{N}$, omitting any identity elements $b_{i}$.
We adopt the following notational conventions (conventions vary in the literature). Strands are enumerated from top to bottom with local indices (i.e. distance from top at that position in the diagram). The diagram is read left to right. The singularities are labeled $T_{i}$ for a tangency between the $i^{th}$ and $i+1^{st}$ strands, and $I^{i}_{i+k}$ for a transverse intersection between the $i^{th}$ through $(i+k)^{th}$ strands. Braid words, read left to right, are in multiplicative notation, $\sigma_{i}$ indicating a crossing between the $i^{th}$ and $(i+1)^{st}$ strands where the $i^{th}$ strand crosses over the $(i+1)^{st}$ strand from left to right. See Figure 6 for an example.
Figure 6. Braided wiring diagram $T_{2}\sigma_{1}^{-1}\sigma_{3}^{-1}T_{2}I^{3}_{4}I^{3}_{4}I^{3}_{4}\sigma_{3}^{-1}I^{2}_{3}\sigma_{2}^{-1}I^{3}_{4}\sigma_{3}\sigma_{2}I^{1}_{2}I^{2}_{4}$, or $(1,T_{2},\sigma_{1}^{-1}\sigma_{3}^{-1},T_{2},1,I^{3}_{4},1,I^{3}_{4},1,I^{3}_{4},\sigma_{3}^{-1},I^{2}_{3},\sigma_{2}^{-1},I^{3}_{4},\sigma_{3}\sigma_{2},I^{1}_{2},1,I^{2}_{4})$ in expanded form.
@@
Fix $0<\mu\ll\pi/2$. Our model for a tangency will be the complex curve $\{(x,y)\in\mathbb{C}^{2}\mid e^{i\mu}x=y^{2}\}$ with the tangency of $\pi_{x}$ at the origin. The braided wiring diagram and braid monodromy near this model are shown in Figure 7. The preimage of the model over the interval in the $Re(x)$ axis includes the singularity, and gives the picture for the braided wiring diagram near a tangency. The braid monodromy about a small circle oriented counterclockwise around the singularity is a single positive half twist on two strands. We break up this circle into two half-circles showing the push-off of the singular braid in the positive and negative imaginary directions in $\mathbb{C}_{x}$. Pushing off in the positive imaginary direction gives the braid $\sigma_{i}^{-1}$ read left to right. Pushing off in the negative imaginary direction gives the trivial braid. The counterclockwise monodromy around the singularity is obtained by composing the inverse of the positive pushoff braid with the negative pushoff braid (since a counterclockwise circle travels right to left along the top and left to right along the bottom). Thus the total monodromy about the tangency is $\sigma_{i}$ as expected.
@@
The model for a transverse intersection is given by a pencil of complex lines which are the complexification of real transverse lines. See [PS23, CS97] for more details on braided wiring diagrams with only transverse intersection singularities and models. Figure 8 shows a braided wiring diagram including both tangency and transverse intersection models, along with the pushoffs in the positive and negative imaginary directions. Let $\Delta_{j,j+k}$ denote the positive half twist on strands $j,j+1,\dots,j+k$ (e.g. $\Delta_{j,j+1}=\sigma_{j}$). Then the positive pushoff of $I^{i}_{i+k}$ is $\Delta_{i,i+k}^{-1}$ and the negative pushoff is $\Delta_{i,i+k}$. Thus, for the surface bundle in the complement of the curve over a small circle around the image of the transverse intersection, the monodromy is $\Delta_{i,i+k}^{2}$, the positive Dehn twist about a curve convexly enclosing holes $i,i+1,\dots,i+k$.
@@
[…]
@@
Let $(\Gamma,\{p_{j}\})$ be a DJVS arrangement with marked points and the corresponding braided wiring diagram $(b_{0},S_{1},b_{1},S_{2},\dots,b_{N-1},S_{N},b_{N})$, where a free marked count will also be considered a singular point $S_{j}$, treated as an intersection point with only one strand. Fix an identification of the planar fiber $P$, a disk with disjoint holes of equal radius centered along a straight line. If the $j^{th}$ singularity is an intersection $I^{i}_{i+k}$, let $A_{j}$ denote a convex curve containing holes $i,i+1,\dots,i+k$, and let $\phi_{j}=\Delta_{i,i+k}^{-1}$ denote the negative half twist. (In the case of a free marked point where $k=0$, $A_{j}$ is the boundary parallel curve around the $i^{th}$ hole, $\phi_{j}:=\Delta_{i,i}^{-1}$ is the negative half twist on $A_{j}$). If the singularity $S_{j}$ is a tangency between strands $i,i+1$, let $A_{j}$ denote a straight line arc connecting hole $i$ and $i+1$, and let $\phi_{j}:=\Delta_{i,i+1}^{-1}=\sigma_{i}^{-1}$ be the negative half twist along $A_{j}$.
Let $(V_{1},V_{2},\dots,V_{N})$ be an ordered collection of vanishing cycles and arcs corresponding to the nearly Lefschetz fibration on the manifold $W_{(\Gamma,p)}$ produced by Theorem 3.8. Then up to generalized Hurwitz equivalence,
@@
Above, $b_{i}\in Map(P)$ stands for an arbitrary lift of the braid $b_{i}$ to the mapping class group of the disk with holes. Different choices yield the same vanishing cycles and may change vanishing arcs by framing conjugation only.
@@
Recall that $W_{(\Gamma,p)}=[(D_{x}\times D_{y})\#_{n}\overline{\mathbb{C}\textup{P}}\,\!^{2}]\setminus\cup\nu(\widetilde{\Gamma}_{i})$, and the Lefschetz fibration is the composition of the blow-down with the projection to $D_{x}$. Let $\alpha:I\to D_{x}$ be the arc over which lies the braided wiring diagram with data $(b_{0},S_{1},b_{1},S_{2},\dots,b_{N-1},S_{N},b_{N})$. We will define vanishing cycles and arcs in the fiber over $\alpha(0)$. We can identify the vanishing cycles by choosing an ordered collection of loops, one about each singular value, such that the composition of the loops is homotopic to an embedded loop convexly surrounding all the singular values. The monodromy about each loop will be a positive (right handed) Dehn twist about the vanishing cycle or a positive boundary interchange about the vanishing arc, depending on whether the singularity is Lefschetz or exotic respectively. Thus if we can choose such a collection of loops in $D_{x}$, determine the monodromy above each loop, and realize this monodromy as a positive Dehn twist about a curve, then we will have identified a set of vanishing cycles for the Lefschetz fibration. The choices made in this process may lead to different collections of vanishing cycles/arcs but any two such sets of data are related by generalized Hurwitz equivalence.
Let $t_{1},\dots,t_{N}\in(0,1)$ be such that $q_{i}:=\alpha(t_{i})$ is the image of the singularity $S_{i}$. Let $t_{0}=0$, $t_{N+1}=1$. Choose a sufficiently small $\delta>0$ such that $\alpha|{[t_{i}-\delta,t_{i}+\delta]}(t)={(t-t_{i})}+q_{i}$ is parallel to the real axis in $D_{x}$ and the braids $b_{i}$ lie over the intervals $\alpha([t_{i}+\delta,t_{i+1}-\delta])$ for $i=0,\dots,N$. Let $h_{i}(t)=\delta e^{\pi i(1-t)}+q_{i}$ be the clockwise path along the upper semi-circle of radius $\delta$ around $q_{i}$. Let $g_{i}(t)=\delta e^{{2\pi}it}+q_{i}$ be the counterclockwise circle of radius $\delta$ around $q_{i}$. Then the loop $l_{i}$ is the concatenation of the following paths: